# Statistical Demography Meets Ministry of Health: The Case of the Family Planning Estimation Tool

### v1

@@ 4.6 Model validation
Model development and checking took place in different stages, based on research on transition models and data models. We carried out a validation exercise in which we fit FPET to a training data set that excluded survey data that had been collected in or after 2018. The goal was to check how well the estimates and predictions based on the training set, i.e. estimates and forecasts that would have been produced in 2018, aligned with: (1) the estimates based on the full data set; and (2) the held-out survey data reported for years after 2018. We calculated summary measures based on prediction errors and calibration.
Detailed findings are in the appendix Section 8.2. In summary, we find that FPET is reasonably well calibrated for predicting mCPR. The point predictions that would have been constructed in 2018 generally fall within the uncertainty intervals constructed using all data at nominal levels or above. Larger absolute errors occur in settings where predictions are off due to an unforeseen stall or acceleration, or where recent data suggests an erroneous trend. In such instances, predictive performance may improve when including EMUs (Mooney et al. [2024a]).
@@ 5 Discussion
In this paper, we described the Family Planning Estimation Tool and the interplay between statistical modeling and usage in practice. We now turn to discussing takeaways and lessons learned from FPET that we believe are relevant to statistical modeling for demographic and global health indicators more generally. We end with some open challenges that also generalize to the broader field.
5.1 Takeaways and lessons learned
5.1.1 Collaborations between users and modelers are a key to success
Based on our experience with statistical modeling for family planning, as well as other reproductive, maternal, and child health indicators, we have concluded that the most important (but often overlooked) modeling challenge is to produce results that are reasonable for all populations and settings. An 80-20 rule in the context of statistical modeling for global health indicators would be as follows: producing reasonable estimates and forecasts for 80% of population-periods is easy; the remaining 20% is hard. While an easy-to-build-low-on-assumptions-type model based on, for example, commonly used generic spatio-temporal smoothing methods, may predict well when considering average results across populations and settings, it typically produces unrealistic results in a subset of settings. Moreover, the 20% of population-periods where models may fail to produce appropriate estimates are often those settings where the estimation matters most, as data are typically sparse or uncertain, or the outlying-ness of outcomes is something that needs to be reflected in the model-based estimates.
To obtain realistic model-based estimates for all settings of interest, including the 20% of population-periods that are unusual in some sense or where data are limited, models need to be built by a collaboration of statisticians and subject experts. Described in a less formal way (see Figure 8): to obtain realistic model-based estimates for all settings of interest, models need to be built by a collaboration between “data crunching” Bugs Bunnies and “this is how the world works” Brainy Smurfs. While the Bugs Bunnies can “easily” introduce powerful statistical tools, the Brainy Smurfs are able to provide reality checks regarding appropriateness of assumptions made and resulting findings. Indeed, the evolution of FPET, in terms of its set-up and assumptions made, was driven by a collaboration of data crunchers and substantive experts through the Track20 project. Active usage and feedback on modeled results resulted in the various updates that we described in earlier sections.
Figure 8: Bugs Bunny, Brainy Smurf, and Bayes. To obtain realistic model-based estimates for all settings of interest, models need to be built by a collaboration between “data crunching” Bugs Bunny and “this is how the world works” Brainy Smurf. The Bayesian framework can be considered for producing model-based findings.
5.1.2 Local tools are another key to success
We believe that local tools are incredibly important for successfully producing locally-relevant estimates. At a minimum, these tools allow for production of estimates based on data from a particular setting alone. It is typically easier to get buy-in for a set of estimates that is produced in-house, based on a fit to data that has been vetted, as opposed to estimates that have been produced in a top-down approach.
@@ 5.1.2 Local tools are another key to success
But the use and usefulness of local models go well beyond the production of just one set of estimates. Having a model that can be run annually builds a mechanism for data review, of surveys but most importantly service statistics. Track20’s experience suggests that countries have improved service statistics over the project because they are reviewing them annually, usually in a consensus meeting with many stakeholders, and it prompts follow-up action. Local tools also enable sensitivity analyses of various forms: models can be fitted to specific data sets or based on changing model assumptions. These kinds of fits can be helpful for analysts to better understand the data and model. Track20 has successfully used this approach in consensus meetings, as well as in additional technical meetings with various government and non-government stakeholders. For example, in India and Kenya, the tool has been used to analyze data outliers and inform decisions related to data exclusion.
In summary, by enabling in-house production of estimates and additional assessments, local tools empower governments and move away from a commonly used approach of externally produced global health estimates.
5.1.3 Communication and training needs to be geared at different audiences
Given the range of audiences involved in modeling exercises, clearly communication needs to be geared at these different audiences and, depending on the audience, vary in terms of technical detail. To make this possible, communication of model assumptions and explanations of “how the model works” are best produced by different actors as well. In our case, the statisticians produced technical writings while Track20 focused on producing materials and supportive documentation that was targeted to less statistical audiences. The local tool can also be helpful for communication, e.g. through sensitivity analyses that illustrate how the model responds when fitted to different sets of data.
5.1.4 Go Bayesian!
@@
We have found that the use of Bayesian models enabled the development and evolution of FPET. The Bayesian framework provides flexibility in translating assumptions into probabilistic models. The wide availability of Bayesian modeling software facilitates the easy implementation of such models and active development of computational approaches have resulted in reduced computation times and expanded the types of models that can be fitted. Last but not least, the Bayesian paradigm also allows for uncertainty to be accounted for in, and propagated across various model components. In our set-up, we derived a local tool from a global model, as a pragmatic, solution-focused approach. The Bayesian framework facilitates this type of set up.
5.2 What’s next: Open challenges
5.2.1 How to allow for context-specific adjustments to model-based estimates?
The use of model-based estimates can be a double-edged sword: On the one hand, model-based estimates can provide insights but on the other hand, model-based estimates may be based on assumptions that do not hold true or estimates may be subject to substantial levels of uncertainty, limiting their usefulness for action. Moreover, as mentioned in relation to the 80-20 rule, the population-periods where models may fail to produce appropriate estimates are often those settings where the estimation matters most, as data are typically sparse or uncertain, or the outlying-ness of outcomes is something that needs to be reflected in the model-based estimates. It is worth noting that the increased focus on capturing within-country differentials compounds these issues: the interest in smaller subgroups brings about greater data needs and – when data availability does not match these needs – exacerbates the importance as well as the potential limitations of models.
When dealing with a setting where model-based estimates conflict with substantive knowledge, i.e. when model-based estimates do not reflect the truth, what can a user or modeler do? Examples from FPET fits are settings where either estimates smooth over a data point (the model considers it an outlier) while expert knowledge indicates that the data point captures the situation very well, or vice versa, where data of dubious quality or from populations that may differ from the population of interest have a great influence on the estimates. In earlier work, we came up with some approaches to deal with some of these settings. For example, users can indicate whether data points refer to populations that may not be representative of the population of interest, and the model fit takes that into account [Alkema et al., 2013]. In the currently-used FPET data models, users can make updates to the default categorization of observations into the “possible outlying” category, to deal with instances where substantive knowledge conflicts with automated categorization of observations into being outlying or not. These FPET approaches are examples of set-ups where users can provide input to improve upon estimates.
While giving the user options to provide additional context related to data or the outcome of interest is a positive way forward to producing high-quality estimates, the open challenge relates to what information to ask for, and how that information can be used. While some context-specific questions may have reasonably clear answers that can be used in modeling, others may not. For example, when considering data quality and errors, it is likely easier to answer the question whether a survey is representative of the entire country or limited to specific regions as opposed to a question related to data quality more generally. In addition to difficulties obtaining relevant information, another challenge is related to the need for standardization of estimation procedures and keeping this process as objective as possible. The option of user inputs should not enable manual tweaking of estimates to optimize gains that are not statistical accuracy but instead, could relate to political agendas. Given these challenges, the process of obtaining context-specific information (and documentation) and its usage needs to be carefully considered.
5.2.2 How to improve the process of model development and communication?
Model development and communication of strengths and weaknesses of model-based estimates is an area of great importance. We believe that the community of statistical modelers have a key role to play in improving on the status quo. The current literature on statistical modeling of global health indicators lacks standardization of the specification and communication of model assumptions. Because of this gap, we find that models are difficult to compare and modeling assumptions may be difficult to understand. In addition, guidance to systematize model building is very limited. This may result in model choice being based on prior experiences of the research team involved as opposed to suitability considerations.
To advance the development and use of model-based estimates, we see a need to develop material to help standardize the specification and communication of model assumptions. To work towards that goal, we proposed the use of a new model class, referred to as Temporal Models for Multiple Populations (TMMP, see Susmann et al. [2022]). This model class facilitates documentation of model assumptions in a standardized way and facilitates comparison across models. We show how existing models for a variety of indicators can be written as TMMPs and how the TMMP-based description can be used to compare and contrast model assumptions. This approach, or other efforts, can be considered to help communicate model assumptions.
There are opportunities to move toward systematizing model building. Given the typical challenges associated with the 80-20 rule, i.e., getting the 20% right, it does not seem realistic to aim for a standard recipe for model building that can be applied across a range of indicators and settings and research questions. However, we do see an opportunity to consider the process of model building in a more systematic manner, through ideas that have been recently put forward under the Bayesian workflow umbrella. In our recent work on FPET, we have introduced elements of this type of workflow that are relevant to our modeling goals. Last but not least, modelers cannot (and should not aim to) pretend that model-based estimates have the final word in settings where additional data are needed to answer open questions. A more targeted approach to data collection, one that is informed by statistical models, could result in great gains in insights.
5.2.3 How to increase diversity in the modeling community?
To maximize the value of statistical models for demographic monitoring, action is needed to expand and diversify the community of actors involved, ranging from end users of model outputs to advanced users (applied statisticians and data scientists who understand the detailed workings of models), and modelers who develop and update models.
To date, the great majority of major modeling efforts are based on a set up where academics or international organizations based in the global north develop models and often also produce estimates. The Track20 project with FPET can serve as an example of how to diversify this community, through local capacity building that enabled local usage of tools and local production and ownership of estimates, and through improved interaction between local users and other actors involved. Additional opportunities are to diversify communities of modelers and advanced users so that there are more statistical models being produced in the global south. The expansion of the community of advanced users is especially useful when thinking about rolling out tools that have more advanced usage options.
The opportunities to diversify communities of modelers and advanced users are contingent upon outreach to, and capacity building of under-represented groups, in particularly based on geography. The good news is that with the availability of high-quality open-source software tools such as R and Stan, and the increased availability and quality of online training material in statistics and data science in various forms, access and training are no longer limited to audiences that are geographically close.
With the combination of initiatives like Track20, that diversify the community involved in model-based estimates on specific areas, and the opportunities provided through open-source software and expanded communities of individuals trained in statistics and data science, a more diverse modeling community could become reality in the short-term.
6 Conclusions
Using our experience with the Family Planning Estimation Tool, we discussed take aways and lessons learned related to the field of statistical modeling for monitoring for demographic and global health indicators. We discussed that (1) Interaction between modelers and users/local experts is essential to producing high quality model-based estimates; (2) Local tools are a key to success; (3) Communication and training needs to be geared at different audiences; and (4) Going Bayesian allows for modeling flexibility. We ended with a more general call to action to the community of statisticians and statistical demographers in the form of steps to consider, to optimize the relevance of statistical modeling in practice. Our call to action focuses on open challenges: (1) How to allow for context-specific model adjustments, for improving estimates in specific settings, (2) How to improve the process of model development and communication, and (3) How to diversify the modeling community. While challenges remain, so do continued advancements in the field of statistics and data science, improvements in access and training opportunities, and the need for high quality measurements. Given all that, we are very optimistic about the future of field of statistical modeling for monitoring for demographic and global health indicators.

### v2

@@ 4.2 Model validation
Model development and checking took place in different stages. To summarize final performance, we carried out a validation exercise in which we fit FPET to a training data set that excluded survey data that had been collected in or after 2020. The goal was to check how well the estimates and predictions based on the training set, i.e. estimates and forecasts that would have been produced in 2020, aligned with: (1) the estimates based on the full data set; and (2) the held-out survey data reported for years after 2019. We calculated summary measures based on prediction errors and calibration.
The first set of measures is focused on the comparison between point estimates and credible intervals (CIs) from the training and full data set. We calculate prediction errors in the FPET point predictions, referring to the difference between the FPET point prediction for a given year and its updated estimate. Here the year 2024 is chosen, to reflect a forecast horizon of 5 years based on the training data that excludes data in and after 2020. We also check how often the point prediction falls outside of the CIs based on all data. We expect these errors to be small and for point predictions to fall inside updated intervals at at least the nominal level. Results are given in Table 1 and suggest that FPET is reasonably well calibrated. For example, the median absolute error in mCPR among all women is 1.7%. Similarly, the point predictions that would have been constructed in 2020 generally fall within the credible intervals constructed using all data at nominal levels or above, see Table 2. We do note a negative bias of -1.8% for predicting mCPR among married women (meaning that FPET overpredicted the mCPR levels). This type of bias was also present in the previously-used FPET model, the updated model has improved the bias from -2.6 to -1.8% (see appendix Section 7.4).
indicator married? MedE MedAE prop <1<1% prop <5<5%
mCPR all -1.1 1.7 0.4 0.8
mCPR no -0.4 1.4 0.4 0.9
mCPR yes -1.8 2.4 0.3 0.7
unmet modern all 0.5 1.2 0.5 1.0
unmet modern no -0.1 0.6 0.7 1.0
unmet modern yes 1.1 1.6 0.3 0.9
Table 1: Summary of prediction errors (in percentage) for the year 2024. An error refers to the difference between the estimate for the indicator based on the full data set and the indicator based on the training data, for the year 2024 (5 year forecast horizon). A positive (negative) error indicates that the prediction based on data up to 2018 underpredicted (overpredicted) the target value. Table columns: MedE = median error, MedAE = median absolute error, prop << 1 or 5 % = proportion of errors less than 1 or 5 percentage points.
indicator marital_status below 90% CI above 90% CI
mCPR married 8.3 1.4
mCPR unmarried 5.6 1.4
unmet married 0.0 1.4
unmet unmarried 1.4 0.0
Table 2: Percentage of estimates for 2024 based on data prior to $20$ that fall outside their respective credible intervals based on all data.. Table columns: below or above 90% CI = percentage of estimates that fall below or above the 90% CI.
The second set of measures is focused on the differences between left-out survey data and point predictions. We do not necessarily expect these differences to be small for all survey data: we only expect small differences for survey data that is subject to small sampling and non-sampling errors. Results for observations from DHS surveys in sub-Saharan Africa, where such errors are generally small, are shown in Table 3. We see that results are comparable to the results in Table 1.
indicator married? MedE MedAE prop <1<1% prop <5<5%
mCPR no -0.5 2.4 0.1 0.8
mCPR yes -1.7 2.1 0.2 0.7
unmet modern no 0.1 0.9 0.6 1.0
unmet modern yes 1.3 2.1 0.3 0.9
Table 3: Summary of prediction errors (in percentage) for left-out DHSs in sub-Saharan Africa. An error refers to the difference between a left-out DHS observation and the estimate for the indicator based on the validation data. A positive (negative) error indicates that the prediction based on data up to but excluding 2020 underpredicted (overpredicted) the target value. Table columns: MedE = median error, MedAE = median absolute error, prop << 1 or 5 % = proportion of errors less than 1 or 5 percentage points.
4.3 Including service statistics data
The effect of including EMUs is illustrated in a case study in Figure 10111EMU data are country-owned and not publicly available. We illustrate EMU data for an anonimized population.. The top row shows information on mCPR (left) and rates of change in mCPR (right), based on EMUs (black) and survey-based estimates of mCPR (red). The black vertical lines indicate the 95% confidence intervals for mCPR based on individual EMUs and their respective $s_{EMU,d,c,t}$ alone. The EMU-based rates of change suggest an acceleration after the most recent survey year, with rates of change that are higher than survey-based predictions. Plots in the bottom row illustrate the effect of including the EMUs in FPET estimates: the high rate of growth in EMUs past the most recent survey results in faster estimated mCPR growth, as compared to survey-only estimates.
mooney_emu assess whether the inclusion of EMUs improves predictions for years after the most recent year with survey data. They find that EMUs improve predictive performance. The inclusion of EMUs is found to be particularly helpful in settings with stalls or accelerations that occurred after the most recent survey year.
Figure 10: Illustration of the effect of including EMUs when estimating mCPR. The top row shows observed EMUs (black) for the level (left) and rate of change (right). The vertical lines indicate the 95% confidence intervals based on $s_{EMU,d,c,t}$ alone. Estimates of mCPR that are based on survey data alone are added in red. The purple dashed line indicates the most recent survey year. The top row illustrates that the EMU-based rates of change suggest an acceleration after the most recent survey year, with rates of change that are higher than survey-based predictions. Plots in bottom row illustrate the effect of including the EMUs in FPET estimates: the acceleration in EMUs past the most recent survey results in faster mCPR growth, as compared to survey-only estimates.
@@ 5 Discussion
In this paper, we introduced the Family Planning Estimation Tool. We summarized the Bayesian hierarchical transition modeling approach underlying FPET and introduced recent updates related to capturing contraceptive transitions, fitting to survey data that may be error prone, and the use of service statistics data. A validation exercise suggested that FPET is reasonably well calibrated for predicting FP outcomes. We presented country-level examples to illustrate FPET’s main utilities: (1) reconciling data from multiple sources with varying levels of uncertainty, and (2) producing backcasts, nowcasts, and forecasts in data-limited settings.
@@
Based on our experience with FPET, we believe that local tools are a key to success for producing locally-relevant estimates. In our set-up, we derived a local tool from a global model, to allow users to fit the model to their own data. At a minimum, these tools allow for production of estimates based on data from a particular setting alone. It is typically easier to get buy-in for a set of estimates that is produced in-house, based on a fit to data that has been vetted, as opposed to estimates that have been produced in a top-down approach. But the use and usefulness of local models go well beyond the production of just one set of estimates. Having a model that can be run annually builds a mechanism for data review, of surveys but most importantly service statistics. Track20’s experience suggests that countries have improved service statistics over the project because they are reviewing them annually, usually in a consensus meeting with many stakeholders, and it prompts follow-up action. Local tools also enable sensitivity analyses of various forms: models can be fitted to specific data sets or based on changing model assumptions. These kinds of fits can be helpful for analysts to better understand the data and model. Track20 has successfully used this approach in consensus meetings, as well as in additional technical meetings with various government and non-government stakeholders. For example, in India and Kenya, the tool has been used to analyze data outliers and inform decisions related to data exclusion. In summary, by enabling in-house production of estimates and additional assessments, local tools empower governments and move away from a commonly used approach of externally produced global health estimates.
@@
The use of Bayesian models enabled the development and evolution of FPET. The Bayesian framework provides flexibility in translating assumptions into probabilistic models and allows for uncertainty to be accounted for, and propagated across various model components. Active development of computational approaches and Bayesian modeling software, such as those by the Stan team, have resulted in reduced computation times and expanded the types of models that can be fitted. More generally, these developments, as well as the sharing of open-source code bases for Bayesian modeling in various initiatives are greatly increasing the availability of Bayesian models for a range of demographic and global health indicators.
We conclude by commenting on the role of the DHS program in FP monitoring and the implications of its recent termination. The DHS has been a cornerstone of FP monitoring in many countries, providing household survey data that inform FPET estimates and a wide range of other analyses (see, for example, [unpd_dhs2025]). Following the DHS program’s recent termination as part of the suspension/termination of several USAID‑funded programs under the new U.S. administration (e.g., see [khaki_when_2025]), the need to combine data from multiple sources and to develop new data systems is more important than ever. In a data landscape that excludes DHS, available sources include surveys conducted by other programs (e.g., MICS, PMA), country‑level surveys, and routine service‑statistics systems. Although each of these sources may have limitations, as discussed in this paper, approaches such as FPET offer a standardized framework for reconciling heterogeneous data and producing coherent FP estimates. Sustained investment in alternative data sources, transparent methods, and country capacity will be critical to maintain robust FP monitoring in the absence of DHS.